OpenStudy (anonymous):

NASA wishes to send a 33 kg satellite into orbit 30,000 km above the SURFACE of the Earth. Calculate the gravitational field at that altitude.

OpenStudy (anonymous):

G=6.67E-11 Mass of earth = 5.98*10^24 radius of earth =6.37*10^6

OpenStudy (anonymous):

The gravitational field \(g\) a distance \(R\) from the center of mass of an object of mass \(m\) can be derived from Newton's law of gravitation. The relationship you'll want is \[g = \frac{GM}{r^2}.\]

OpenStudy (anonymous):

Of course, I meant\[g = \frac{GM}{R^2}\]

OpenStudy (anonymous):

gravitational field = gravitational force /mass of the object therefore g=Gme/r2 where r is the distance from the surfaceof the earth+the radius of the earth